OpenStudy (anonymous):

9y^2 - 12y + 4 = 0 how do you solve that

OpenStudy (anonymous):

(3y-2)(3y-2)=0 or (3y-2)^2=0 which means we set : 3y-2=0 and solve for y.

OpenStudy (anonymous):

what it doesnt make sense

OpenStudy (anonymous):

dont you have to use the formula or do you have to factor it

OpenStudy (anonymous):

you have to factor :9y^2 - 12y + 4 to get the roots or x values. When you factor this you get (3y-2)(3y-2). Now you set these terms to zero and solve for y. That means: 3y-2=0 and 3y-2=0 For both of these, we can add 2 to both sides and get: 3y=2, then we divide both sides by 3 and get: y=2/3

OpenStudy (anonymous):

so does that mean that y = 2/3 but arent there supposed to be two answers

OpenStudy (anonymous):

you see what happened when i factored: 9y^2 - 12y + 4? I got (3y-2)(3y-2). These are the same terms. When i solve both terms for y i get y=2/3. Thus there is only one root y=2/3

OpenStudy (anonymous):

ohh becuase both the factors are the same

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

you could say y = 2/3 D.R. for double root

OpenStudy (anonymous):

what is double root

OpenStudy (anonymous):

both the factors are the same and equal. so it's called a double root..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer would be y = 2/3

OpenStudy (anonymous):

yes the answer would be y=2/3

OpenStudy (anonymous):

and can you help me with another one